A person riding on a Ferris Wheel of radius 14.0 m at some moment has a tangential speed of 2.00 m/s (rotating clockwise) and is slowing down at a rate of 1.00x10-1 m/s2 . At the middle point on the right, indicated by the red circle, determine the magnitude and exact direction of his acceleration vector.

140m
Answer #1

Radius of the wheel r = 14 m/s

Tangential acceleration is aT = -0.10 m/s2, Centripetal acceleration is given by

a_{c}= \frac{v^{2}}{r}= \frac{(2 m/s)^{2}}{14 m} = 0.286 m/s^{2}

Net magnitude of the acceleration is

a =\sqrt{ a_{c}^{2}+a_{T}^{2}}=\sqrt{ (0.286 m/s^{2})^{2}+(-0.10 m/s^{2})^{2}}= 0.303 m/s^{2}

Direction will be

\Phi = \tan^{-1}\left ( \frac{a_{T}}{a_{c}} \right )= \tan^{-1}\left ( \frac{-0.100}{0.286} \right )= -19.3^{0}

Negative sign indicates clocwise angle wrt radius vector or centripetal acceleration vector.